An electronic company in Laguna manufactures resistors that have a mean resistance of 120 ohms and a standard deviation of 13 ohms. Find the probability that a random sample of 40 resistors will an average resistance greater than 114 ohms.

1
Expert's answer
2022-03-14T17:11:31-0400

"P(\bar X>114)=P(Z>\frac{114-120}{\frac{13}{\sqrt{40}}})=P(Z>-2.92)=1-P(Z<-2.92)=0.9982."